# DSP: Dynamic Sequence Parallelism for Multi-Dimensional Transformers

### v2

@@ Abstract
Scaling multi-dimensional transformers to long sequences is indispensable across various domains. However, the challenges of large memory requirements and slow speeds of such sequences necessitate sequence parallelism. All existing approaches fall under the category of embedded sequence parallelism, which are limited to shard along a single sequence dimension, thereby introducing significant communication overhead. However, the nature of multi-dimensional transformers involves independent calculations across multiple sequence dimensions. To this end, we propose Dynamic Sequence Parallelism (DSP) as a novel abstraction of sequence parallelism. DSP dynamically switches the parallel dimension among all sequences according to the computation stage with efficient resharding strategy. DSP offers significant reductions in communication costs, adaptability across modules, and ease of implementation with minimal constraints. Experimental evaluations demonstrate DSP’s superiority over state-of-the-art embedded sequence parallelism methods by remarkable throughput improvements ranging from 32.2% to 10x, with less than 25% communication volume.
@@ 1 Introduction
Efficiently scaling multi-dimensional transformers to accommodate long sequences is necessary across diverse domains, including video generation [2, 29, 45], image generation [27, 38, 41], protein structure prediction [21], spatial-temporal information processing [6], and beyond. The long length of sequences in these applications entails substantial activation memory costs and a notable slowdown in processing speeds, underscoring the need for employing sequence parallelism.
Current methods for sequence parallelism, such as Megatron-SP [22], DeepSpeed-Ulysses [20], and Ring-Attention [25, 26] are all embedded sequence parallelism methods. As shown in Figure 1, these embedded methods shard along a single sequence dimension, which are tailored to the task-specific pattern of modules and introduce intricate communication and complex code modification. However, multi-dimensional neural networks often operate independent computations across multiple sequence dimensions. For instance, for video generation models like OpenSora [50] and Latte [29], Spatial-Temporal Attention [48] is favored over full Self-Attention [46], facilitating separate attention computations for temporal and spatial dimensions.
@@
Therefore, there exists an explorable space for a general sequence parallelism abstraction. To adapt to the flexible patterns of multi-dimensional transformers, we introduce Dynamic Sequence Parallelism (DSP) as a novel abstraction of sequence parallelism, characterized by its elegance, high effectiveness, and excellent compatibility with popular transformers. Unlike embedded sequence parallelism, DSP dynamically switches the parallel dimension of sequences during the computation stage with an efficient resharding strategy, a process completely decoupled from the computation module. DSP offers several advantages over embedded sequence parallelism: 1) Efficient Communication: DSP incurs significantly lower communication costs due to its simplified communication patterns and reduced frequency of exchanges. 2) Adaptability: DSP seamlessly adapts to most modules without necessitating specific modifications and imposes few limitations on its usage. 3) Ease of Use: DSP is remarkably easy to implement, and also provides a simple API shown in Appendix A.4 for users to enable it effortlessly.
Our experiments yield promising results, showcasing DSP’s superiority over state-of-the-art embedded sequence parallelism methods. It achieves an end-to-end throughput improvement ranging from 32.2% to 10x and reduces communication volume by at least 75%.
@@
  • By significantly reducing communication volume and frequency, DSP improves end-to-end throughput by 32.2% to 10x and reduces communication volume by at least 75% compared to state-of-the-art methods.
@@ 2 Background and Related Work
$B$ The number of batch sizes $N$ The number of GPUs
$C$ The size of hidden states $n$ The $n$-th GPU
$S_{i}$ The $i$-th sequence dimension $\mathbf{X}$ The a multi-dimensional sequence
$s_{i}$ The status that sequence is sharded from tensor
dimension $S_{i}$ $\mathbf{X}_{p,n}$ The partition of X assigned to GPU $n$
$\hat{s}$ The status that sequence is not sharded for sequence parallel
$M$ The volume of a sequence tensor
@@ 2.1 Background
Transformer Architecture. Transformer [46] is a type of neural network architecture that has become highly influential in natural language processing [3, 9, 40] and other domains [10, 34, 21]. The Transformer is composed of a stack of layers, each consisting of a multi-head attention (MHA) and a position-wise feed-forward network (FFN). Specifically, the MHA comprises H independently parameterized attention heads, formulated as:
$\mathrm{MHA}(x)=\mathrm{Concat}(\mathrm{head}_{1},\dots,\mathrm{head}_{H})\mathbf{W}^{O},\quad\mathrm{head}_{i}=\mathrm{Att}(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}),$ (1)
$\mathrm{Att}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},\quad x_{\mathrm{MHA}}=\mathrm{LayerNorm}(x+\mathrm{MHA}(x)),$ (2)
@@
$\mathrm{FFN}(x)=\max(0,x\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}_{2},\quad x_{\mathrm{out}}=\mathrm{LayerNorm}(x_{\mathrm{MHA}}+\mathrm{FFN}(x_{\mathrm{MHA}})),$ (3)
@@
Figure 2: Illustration of Multi-Dimensional (2D in this case) Transformer.
Multi-Dimensional Transformer. Multi-dimensional transformers [17, 49] extend the self-attention mechanism of standard transformers to operate over multiple dimensions beyond just one sequence. An example of 2D-Transformer is shown in Figure 2. Let the input multi-dimensional sequence be denoted as $\mathbf{X}\in\mathbb{R}^{[B,S_{1},S_{2},\dots,S_{K},C]}$, where $B$ is the batch size, $S_{1},S_{2},\dots,S_{K}$ are the sequence lengths along $K$ different sequence dimensions, and $C$ is the hidden size. Multi-dimensional transformer can be formatted as:
@@
After applying the transformer block operation along all $N$ dimensions, the final output tensor $\mathbf{X}_{\mathrm{out}}$ has the same shape as the input tensor $\mathbf{X}$. Multi-dimensional Transformer is widely used for applications with multi-dimensional inputs including video data [29, 47, 14, 11], 3D data [52, 4], protein structure prediction [21, 30], time-series data [32, 8, 18] and beyond.
@@ 2.2 Related Work
Data parallelism [16, 24] is one of the most widely adopted parallelism techniques. The input data is partitioned across devices, each processing a subset. Model parameters are replicated, and gradients are summed. ZeRO [36, 37] optimizes memory by partitioning parameters, states, and gradients across devices, enabling the training of larger models. Tensor parallelism [44, 42], or model parallelism, partitions model parameters across devices. Different model parts are assigned to different devices. Pipeline parallelism [19, 31, 23, 28] partitions the model into stages executed in parallel across devices. Activations are passed between devices in a pipeline style.
@@
Ring Attention. Li et al. [25] employs an innovative approach to partitioning the sequence dimension using a ring-style peer-to-peer (P2P) communication pattern to transfer keys and values across GPUs. Liu et al. [26] enhance this method with an online softmax mechanism, allowing for the computation of attention scores without retaining the full sequence length. However, Ring Attention’s reliance on P2P communication can be less efficient in high-latency environments.
@@
Megatron-SP. Korthikanti et al. [22] further optimizes activation usage in the attention based on tensor parallelism. To transit between tensor parallelism and sequence parallelism in the Transformer block, additional all-gather and reduce-scatter operations are introduced. And it’s constrained by the number of attention heads, as self-attention relies on the parallelism of the head dimension of the sequence.
@@
DeepSpeed-Ulysses. Jacobs et al. [20] introduces an innovative approach for training long sequences by utilizing all-to-all collective communication. This method partitions the query, key, and value matrices across attention heads while preserving the original attention computation structure. The process is facilitated by two sets of all-to-all communications that alternate between sequence splitting and attention head splitting. Nevertheless, it is constrained by the number of attention heads as well.
@@
Moreover, these sequence parallelism methods are designed for parallelism within a single sequence dimension. For multi-dimensional transformers, this strategy becomes inefficient due to unnecessary communication for the head dimension. While specialized parallelism for long sequences and complex operations has been explored in specific domains [5], their applicability remains limited.
3 Dynamic Sequence Parallelism
3.1 Overview
Figure 3: System overview of Dynamic Sequence Parallelism.
In the realm of multi-dimensional transformers, computation occurs independently for each sequence dimension. To harness this inherent feature, we introduce Dynamic Sequence Parallelism (DSP), an efficient, adaptive and ease-of-use sequence parallelism abstraction for multi-dimensional transformers.
To ensure correct computation logic with sequence parallelism, embedded methods typically require complex and time-consuming communications within computation modules to change the parallel dimension. However, as illustrated in Figure 3, the key feature of DSP is its dynamic switch of parallel dimension at the intersections of computation stages. By resharding only between computation stages dynamically, rather than within them, this approach allows DSP to remain independent of the computation logic within the module. Therefore, DSP eliminates numerous unnecessary communications within modules, and is able to utilize efficient all-to-all operations to switch parallelism dimensions for the intermediate sequence.
For operations involving all sequence dimensions, including the beginning and end of the model, DSP handles them by split and gather operation. Furthermore, we also propose a high-level, user-friendly implementation of DSP compatible with all distributed frameworks based on PyTorch, details in Appendix A.4.
3.2 Problem Definition
Given a multi-dimensional sequence $\mathbf{X}\in\mathbb{R}^{[B,S_{1},S_{2},\dots,S_{K},C]}$ and a set of $N$ GPUs, where $S_{1},\dots,S_{K}$ are the sequence along $K$ different sequence dimensions, we aim to partition the computation such that the memory usage per GPU is under capacity while maintaining acceptable communication costs. Let $\mathbf{X}_{p,n}$ denote the partition of $\mathbf{X}$ assigned to GPU $n$, where $p$ represents the partition strategy. The optimization problem is formulated as:
$\min_{p}\sum_{n=1}^{N}\mathrm{CommCost}(\mathbf{X}_{p,n})\ \ s.t.\ \ \mathrm{Memory}(\mathbf{X}_{p,n})<Capacity.$ (6)
Here, $\mathrm{Memory}(\mathbf{X}{p,n})$ denotes the memory usage of partition $\mathbf{X}_{p,n}$ on GPU $n$, $\mathrm{CommCost}(\mathbf{X}_{p,n})$ represents the communication cost. We aim to achieve a balance that minimizes the overall computational overhead while optimizing GPU resource utilization.
3.3 Dynamic Primitives
@@
Figure 4: Illustration of dynamic switch.
In this section, we introduce the key dynamic primitives of DSP, as outlined in Table 2. These three primitives form the cornerstone for implementing DSP across a spectrum of multi-dimensional transformers.
@@
where the resulting tensor $\mathbf{Y}$ has the shape $\mathbb{R}^{[B,S_{1},\dots,S_{i},\dots,S_{j}/N,\dots,S_{K},C]}$. Furthermore, Split and Gather operations facilitate smooth transitions between sharded and non-sharded states. Although these operations may involve increased communication compared to Switch operations, they are primarily utilized at the onset and conclusion of most networks, and also for some global operations in very rare conditions, rendering their costs negligible.
@@ 3.4 Adaptability and Flexibility
Given its decoupling from the computation of modules, DSP exhibits remarkable adaptability, making it compatible with a wide array of transformer variants such as Cross Attention [29, 15]; specialized kernels like FlashAttention [7]; special attention mechanisms including multi-query attention [43] and grouped-query attention [1]; and even beyond like Mamba [12] and RWKV [35]. This inherent flexibility enables DSP to seamlessly integrate into diverse transformers without specific modification. Furthermore, while DeepSpeed-Ulysses and Megatron-LM necessitate attention head splitting, DSP’s scalability is significantly better because it shards on sequence length, which much larger especially when scaling sequences.
@@
As shown in Appendix A.4, we demonstrate the API usage of DSP. By calling just four functions without knowing the detailed implementation, DSP can be enabled on PyTorch and is compatible with various distributed frameworks, including FSDP [51], Accelerate [13], DeepSpeed [39], and Megatron-LM [44].

### v5

@@ Abstract
Scaling multi-dimensional transformers to long sequences is important across various domains. The challenges of large memory requirements and slow speed of such sequences require sequence parallelism. All existing approaches fall under the category of embedded sequence parallelism, which are limited to shard along a single sequence dimension, thereby introducing significant communication overhead. However, multi-dimensional transformers involve independent calculation across multiple sequence dimensions. To this end, we propose Dynamic Sequence Parallelism (DSP) as a novel abstraction of sequence parallelism. DSP dynamically switches the parallel dimension according to the computation stage with efficient resharding strategy. DSP offers significant reductions in communication costs, adaptability across modules, and ease of use with minimal constraints. Experiments demonstrate DSP’s superiority over state-of-the-art sequence parallelism methods by remarkable throughput improvements ranging from 32.2% to 10$\times$, with at least 50% communication volume reduction.
Machine Learning, ICML
@@ 1 Introduction
Efficiently scaling multi-dimensional transformers to accommodate long sequences is necessary across diverse domains, including video generation (Singer et al., 2022; Blattmann et al., 2023; Ma et al., 2024), image generation (Ramesh et al., 2021; Rombach et al., 2022; Liu et al., 2024), protein structure prediction (Jumper et al., 2021), spatial-temporal information processing (Cong et al., 2021), and beyond. The long length of sequences introduces substantial activation memory costs and notable slowdown for speed, underscoring the need for employing parallelism.
Apart from data parallel and pipeline parallel (Huang et al., 2019) which cannot reduce memory cost and inference time, sequence parallel is the only option. Current sequence parallelism, such as Megatron-LM (Shoeybi et al., 2019), Ring-Attention (Li et al., 2021; Liu et al., 2023a), Megatron-SP (Korthikanti et al., 2022), and DeepSpeed-Ulysses (Jacobs et al., 2023) are all embedded sequence parallelism methods. As shown in Figure 1, these embedded methods shard along a single sequence dimension, which are tailored to the specific pattern and introduce extra communication and complex code modification.
However, multi-dimensional transformers calculate independently across multiple sequence dimensions. For instance, for video generation models like OpenSora (Zangwei Zheng, 2024) and Latte (Ma et al., 2024), Spatial-Temporal Attention (Yan et al., 2021) is adopted which separates attention computations to independent temporal and spatial computation. Therefore, there exists a potential space for a new sequence parallelism paradigm.
@@
To adapt to the flexible patterns of multi-dimensional transformers, we introduce Dynamic Sequence Parallelism (DSP) as a novel abstraction of sequence parallelism, featured by its elegant design, high effectiveness, and excellent compatibility. Unlike embedded sequence parallelism, DSP dynamically switches the parallel dimension of sequences during the computation stage with an efficient resharding strategy, completely decoupled from the modules’ logic.
DSP offers several advantages over embedded sequence parallelism: 1) Efficient communication: DSP incurs significantly lower communication costs due to its simplified communication patterns and reduced frequency of exchanges. 2) Adaptability: DSP seamlessly adapts to most modules without necessitating specific modifications and imposes few limitations on its usage. 3) Ease of use: DSP is remarkably easy to implement, and also provides a simple API for users to enable it effortlessly.
Our experiments yield promising results, showcasing DSP’s superiority over state-of-the-art embedded sequence parallelism methods. It achieves an end-to-end throughput improvement ranging from 32.2% to 10$\times$ and reduces communication volume by at least 75%.
@@
  • By significantly reducing communication volume and frequency, DSP improves end-to-end throughput by 32.2% to 10$\times$ and reduces communication volume by at least 50% compared to state-of-the-art methods.
@@ 2 Related Work
$B$ The number of batch sizes.
$C$ The size of hidden states.
$M$ The volume of a sequence tensor.
$N$ The number of GPUs.
$n$ The $n$-th GPU.
$\mathbf{X}$ The a multi-dimensional sequence.
$\mathbf{X}_{p,n}$ The partition of X assigned to GPU $n$.
$S_{i}$ The $i$-th sequence dimension.
$s_{i}$ The status that sequence is sharded from
dimension $S_{i}$.
$\hat{s}$ The status that sequence is not sharded.
@@ 2.1 Background
Transformer Architecture. Transformer (Vaswani et al., 2017) is a type of neural network architecture that has become highly influential in natural language processing (Devlin et al., 2018; Brown et al., 2020; Reid et al., 2024) and other domains (Dosovitskiy et al., 2020; Jumper et al., 2021; Peebles & Xie, 2022). The Transformer is composed of a stack of layers, each consisting of a multi-head attention (MHA) and a position-wise feed-forward network (FFN). Specifically, the MHA comprises H independently parameterized attention heads, formulated as:
$\mathrm{MHA}(x)=\mathrm{Concat}(\mathrm{head}_{1},\dots,\mathrm{head}_{H})\mathbf{W}^{O},$ (1)
$\mathrm{head}_{i}=\mathrm{Att}(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i}).$ (2)
$\mathrm{Att}(\mathbf{Q},\mathbf{K},\mathbf{V})=\mathrm{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V},$ (3)
$x_{\mathrm{MHA}}=\mathrm{LayerNorm}(x+\mathrm{MHA}(x)).$ (4)
@@
$\mathrm{FFN}(x)=\max(0,x\mathbf{W}_{1}+\mathbf{b}_{1})\mathbf{W}_{2}+\mathbf{b}_{2},$ (5)
$x_{\mathrm{out}}=\mathrm{LayerNorm}(x_{\mathrm{MHA}}+\mathrm{FFN}(x_{\mathrm{MHA}})),$ (6)
@@
Figure 2: Illustration of multi-dimensional (2D for this example) transformer. It calculates each dimension of sequence independently at the corresponding stage. After the calculation is done in one dimension, it will switch to another dimension in next stage.
Multi-Dimensional Transformer. Multi-dimensional transformers (Ho et al., 2019; Yang et al., 2022) extend the self-attention mechanism of standard transformers to operate over multiple dimensions beyond just one sequence. An example of 2D-Transformer is shown in Figure 2. Let the input multi-dimensional sequence be denoted as $\mathbf{X}\in\mathbb{R}^{[B,S_{1},S_{2},\dots,S_{K},C]}$, where $B$ is the batch size, $S_{1},S_{2},\dots,S_{K}$ are the sequence lengths along $K$ different sequence dimensions, and $C$ is the hidden size. Multi-dimensional transformer can be formatted as:
@@
After applying the transformer block operation along all $N$ dimensions, the final output tensor $\mathbf{X}_{\mathrm{out}}$ has the same shape as the input tensor $\mathbf{X}$. Multi-dimensional Transformer is widely used for applications with multi-dimensional inputs including video data (Xu et al., 2020; He et al., 2021; Geng et al., 2022; Ma et al., 2024), 3D data (Zheng et al., 2021; Chen et al., 2023), protein structure prediction (Jumper et al., 2021; Mirdita et al., 2022), time-series data (Pan et al., 2022; Huang et al., 2022; Deihim et al., 2023) and beyond.
@@ 2.2 Related Work
Data parallelism (Hillis & Steele Jr, 1986; Li et al., 2020) is one of the most widely adopted parallelism techniques. The input data is partitioned across devices, each processing a subset. Model parameters are replicated, and gradients are summed. ZeRO (Rajbhandari et al., 2019, 2021) optimizes memory by partitioning parameters, states, and gradients across devices, enabling the training of larger models. Tensor parallelism (Shazeer et al., 2018; Shoeybi et al., 2019), or model parallelism, partitions model parameters across devices. Different model parts are assigned to different devices. Pipeline parallelism (Huang et al., 2019; Narayanan et al., 2019; Li & Hoefler, 2021; Liu et al., 2023b) partitions the model into stages executed in parallel across devices. Activations are passed between devices in a pipeline style.
@@
Li et al. (2021) employs an innovative approach to partitioning the sequence dimension using a ring-style peer-to-peer (P2P) communication pattern to transfer keys and values across GPUs. (Liu et al., 2023a) enhance it with an online softmax mechanism, allowing for the computation of attention scores without retaining the full sequence length. However, Ring Attention’s reliance on P2P communication can be less efficient in high-latency environments.
Megatron-LM.
Shoeybi et al. (2019) introduces tensor tensor parallelism by partitioning model across devices. The method splits both feed-forward networks and self-attention layers along their hidden dimension, with necessary all-reduce operations. This approach enables efficient distributed training while reducing memory and communication overhead compared to data parallelism. It also reduces the tensor size on each devices.
Megatron-SP.
@@
Korthikanti et al. (2022) further optimizes activation usage in the attention based on tensor parallelism. To transit between tensor parallelism and sequence parallelism in the transformer block, additional all-gather and reduce-scatter operations are introduced. But it’s limited by the number of attention heads, as self-attention relies on the parallelism of the head dimension of the sequence.
DeepSpeed-Ulysses.
@@
Jacobs et al. (2023) introduces an innovative approach for training long sequences by utilizing all-to-all collective communication. This method partitions the query, key, and value matrices across attention heads while preserving the original attention computation structure. The process is facilitated by two sets of all-to-all communications that alternate between sequence splitting and attention head splitting. Nevertheless, it is constrained by the number of attention heads as well.
@@
Moreover, these sequence parallelism methods are designed for parallelism within a single sequence dimension. For multi-dimensional transformers, this strategy becomes inefficient due to unnecessary communication. While specialized parallelism for multi-dimensional sequences has been explored in specific domains (Cheng et al., 2024), their applicability remains limited.
3 Dynamic Sequence Parallel
@@
Figure 3: Illustration of dynamic switch. Through all-to-all communication, the computation sequence is
3.1 Problem Definition
Given a sequence $\mathbf{X}\in\mathbb{R}^{[B,S_{1},S_{2},\dots,S_{K},C]}$ and a set of $N$ GPUs, where $S_{1},\dots,S_{K}$ are the sequence along $K$ different sequence dimensions, we aim to partition the computation such that the memory usage per GPU is under capacity while maintaining acceptable communication costs. Let $\mathbf{X}_{p,n}$ denote the partition of $\mathbf{X}$ assigned to GPU $n$, where $p$ represents the partition strategy. It can be formulated as:
$\min_{p}\sum_{n=1}^{N}\mathrm{CommCost}(\mathbf{X}_{p,n}),$
$s.t.\ \ \mathrm{Memory}(\mathbf{X}_{p,n})<Capacity.$ (9)
where $\mathrm{Memory}(\mathbf{X}{p,n})$ denotes the memory usage of partition $\mathbf{X}_{p,n}$ on GPU $n$, $\mathrm{CommCost}(\mathbf{X}_{p,n})$ represents the communication cost. We aim to achieve a balance that minimizes the overall computational overhead while optimizing GPU resource utilization.
3.2 Dynamic Primitives
The key dynamic primitives of DSP are outlined in Table 2. These three primitives form the foundation for adapting DSP to various multi-dimensional transformers.
@@ 3.2 Dynamic Primitives
where $\mathbf{Y}$ has the shape $\mathbb{R}^{[B,S_{1},\dots,S_{i},\dots,S_{j}/N,\dots,S_{K},C]}$. Furthermore, Split and Gather operations facilitate smooth transitions between sharded and non-sharded states. Although these operations may involve increased communication compared to Switch operations, they are primarily utilized at the onset and conclusion of most networks, and also for some global operations in very rare conditions, rendering their costs negligible.
3.3 Overview
Figure 4: System overview of Dynamic Sequence Parallelism. It utilizes Split and Gather at the beginning and the end of model to separate and collect the complete sequences. In the middle computation, it utilizes Dynamic Switch to change the sharding dimension with efficient all-to-all communication. So that the following computation will not be affected by sharding.
In the realm of multi-dimensional transformers, computation occurs independently for each sequence dimension. To harness this inherent feature, we introduce Dynamic Sequence Parallelism (DSP), an efficient, adaptive and ease-of-use method for multi-dimensional transformers.
To ensure correct computation logic with sequence parallelism, embedded methods typically require complex and time-consuming communications within computation modules to change the parallel dimension. As illustrated in Figure 4, the key feature of DSP is its dynamic switch of parallel dimension between computation stages.
By resharding only between computation stages dynamically, rather than within them, this approach allows DSP to remain independent of the computation logic within the module. Therefore, DSP eliminates numerous unnecessary communications within modules, and is able to utilize efficient all-to-all operations to switch parallelism dimensions for the intermediate sequence. For operations involving all sequence dimensions, including the beginning and end of the model, DSP handles them by Split and Gather operation.
Furthermore, we also propose a high-level, user-friendly implementation of DSP compatible with all distributed frameworks based on PyTorch.
@@ 3.4 Adaptability and Flexibility
Given its decoupling from the computation of modules, DSP exhibits remarkable adaptability, making it compatible with a wide array of transformer variants such as Cross Attention (Hertz et al., 2022; Ma et al., 2024); specialized kernels like FlashAttention (Dao et al., 2022); special attention mechanisms including multi-query attention (Shazeer, 2019) and grouped-query attention (Ainslie et al., 2023); and even beyond like Mamba (Gu & Dao, 2023) and RWKV (Peng et al., 2023). This inherent flexibility enables DSP to seamlessly integrate into diverse transformers without specific modification. Furthermore, while DeepSpeed-Ulysses and Megatron-SP necessitate attention head splitting, DSP’s scalability is significantly better because it shards on sequence length, which is more redundant.
@@
By calling just four functions without knowing the detailed implementation, DSP can be enabled on PyTorch and is compatible with various distributed frameworks, including FSDP (Zhao et al., 2023), Accelerate (Gugger et al., 2022), DeepSpeed (Rasley et al., 2020), and Megatron-LM (Shoeybi et al., 2019).